# Tight concentration inequalities for quantum adversarial setups exploiting permutation symmetry

### v2

@@ III.2 Application to quantum key distribution
For the next application, we consider Phoenix-Barnett-Chefles 2000 (PBC00) protocol [33], which uses three quantum states forming an equilateral triangle on the $X$-$Z$ plane in the Bloch sphere. The security of this protocol is proved in the finite-size case against general attacks in Ref. [9] using Azuma’s inequality. In order to make the paper as self-contained as possible, we list the protocol below. Let {|$\{\ket{0},\ket{1}\}$ be the $Z$ basis of a qubit. Let {$\{\ket{+},\ket{-}\}$ with |±(|$\ket{\pm}\coloneqq(\ket{0}\pm\ket{1})/\sqrt{2}$ be the $X$ basis. The three states used in PBC00 protocol are |$\ket{\psi_{1}}\coloneqq\cos(2\pi/3)\ket{+}+\sin(2\pi/3)\ket{-}$, |$\ket{\psi_{2}}\coloneqq\cos(4\pi/3)\ket{+}+\sin(4\pi/3)\ket{-}$, and |$\ket{\psi_{3}}\coloneqq\ket{+}$. Let us also define |$\ket{\bar{\psi}_{i}}$ for $i=0,1,2$ that satisfies $\braket{\psi_{i}}{\bar{\psi}_{i}}=0$. Then, the PBC00 protocol is described as follows.
@@
$\begin{split}\mathrm{Pr}\Biggl{[}\frac{\hat{N}_{\rm ph}}{N}-\frac{5}{4}\frac{\hat{N}_{\rm bit}}{N}&\geq\sqrt{\frac{\ln(2/\epsilon)}{2N}}\\ &\quad+\frac{5}{4\sqrt{N}}\biggl{[}\beta-\alpha\Bigl{(}1-\frac{2\hat{N}_{\rm bit}}{N}\Bigr{)}\biggr{]}\Biggr{]}\leq\epsilon,\end{split}$ (189)
@@
Finally, we apply Theorem 3 to this problem. We can regard the measurement in the purified protocol as the three-outcome measurement channel as described above. Since the set of probability mass functions after measurement is included in ${\cal Q}$ in Eq. (193), which is already linear and contains all the extremal points of the convex polytope of the probability mass functions except $(1,0,0)$, we can choose ${\cal R}={\cal Q}$, where ${\cal R}$ is defined in Theorem 3. The region ${\cal R}^{\prime}$ in Theorem 3 is thus given by
@@
IV Discussion
In this paper, we developed concentration inequalities for a permutation-invariant quantum state possibly with an additional symmetry and a sequence of outcomes of independent and identical quantum measurements performed on an adversarial quantum state. For the former case, we used the technique from the representation theory of the permutation group, which is well-known in the community of quantum information theory. With this rather conventional technique, we can upper-bound the permutation-invariant state by i.i.d. quantum states. Compared to the previous de-Finetti-type statement [13, 18, 19], our improvement for this case comes from the fact that one can adaptively choose these i.i.d. states depending on the irreducible subspace of the permutation group. For the latter case, we found an upper bound on the probability of obtaining each sequence of quantum measurement outcomes with a specific type and then obtained a concentration inequality for the probability of obtaining sequences with the atypical types. For this, we developed, to the best of the author’s knowledge, a completely new technique that uses a geometric structure of the convex polytope of the probability mass functions in which the set of types is contained as lattice points. Both bounds we obtained are Sanov-type and are thus tighter than the conventional bound obtained via Azuma’s inequality. In specific models we discussed in Sec. III, our bounds even outperform the recently developed refinement of Azuma’s inequality [12]. Even though the applicability of our bounds is limited compared to Azuma-type inequalities, they still have many applications in QKD as demonstrated in Sec. III.2 and possibly in other quantum information theories to mitigate finite-size effects.
Acknowledgements.

### v3

@@ 3.2 Application to quantum key distribution
For the next application, we consider a QKD protocol with qubits. Even though Bennett-Brassard 1984 protocol [36] may be the most common QKD protocol, its finite-size security analysis can be reduced to a classical random sampling problem [37, 38, 39] in which one can apply a better concentration inequality than Azuma’s and our inequalities. Therefore, we alternatively consider Phoenix-Barnett-Chefles 2000 (PBC00) protocol [40], which uses three quantum states forming an equilateral triangle on the $X$-$Z$ plane in the Bloch sphere. The security of this protocol is proved in the finite-size case against general attacks in Ref. [11] using Azuma’s inequality. In order to make the paper as self-contained as possible, we list the protocol below. Let {|$\{\ket{0},\ket{1}\}$ be the $Z$ basis of a qubit. Let {$\{\ket{+},\ket{-}\}$ with |±(|$\ket{\pm}\coloneqq(\ket{0}\pm\ket{1})/\sqrt{2}$ be the $X$ basis. The three states used in PBC00 protocol are |$\ket{\psi_{1}}\coloneqq\cos(2\pi/3)\ket{+}+\sin(2\pi/3)\ket{-}$, |$\ket{\psi_{2}}\coloneqq\cos(4\pi/3)\ket{+}+\sin(4\pi/3)\ket{-}$, and |$\ket{\psi_{3}}\coloneqq\ket{+}$. Let us also define |$\ket{\bar{\psi}_{i}}$ for $i=0,1,2$ that satisfies $\braket{\psi_{i}}{\bar{\psi}_{i}}=0$. Then, the PBC00 protocol is described as follows.
@@
$\begin{split}&\mathrm{Pr}\Biggl{[}\frac{\hat{N}_{\rm ph}}{N}-\frac{5}{4}\frac{\hat{N}_{\rm bit}}{N}\\ &\qquad\geq\sqrt{\frac{\ln(2/\epsilon)}{2N}}+\frac{5}{4\sqrt{N}}\biggl{[}\beta-\alpha\Bigl{(}1-\frac{2\hat{N}_{\rm bit}}{N}\Bigr{)}\biggr{]}\Biggr{]}\leq\epsilon,\end{split}$ (202)
with
@@
Now we apply Theorem 1 to this protocol. Let ${\cal X}$ be a sample space defined as
@@
and $\bm{p}=(p_{1},p_{2},p_{3})$ be a probability mass function on ${\cal X}$. Then, from Eq. (195), the set ${\cal Q}$ of allowed probability mass function for any quantum state between Alice and Bob can be given by
@@
Finally, we apply Theorem 3 to this problem. We can regard the measurement in the purified protocol as the three-outcome measurement channel as described above. Since the set of probability mass functions after measurement is included in ${\cal Q}$ in Eq. (207), which is already linear and contains all the extremal points of the simplex of probability mass functions except $(1,0,0)$, we can choose ${\cal R}={\cal Q}$, where ${\cal R}$ is defined in Theorem 3. The region ${\cal R}^{\prime}$ in Theorem 3 is thus given by
@@
4 Discussion
In this paper, we developed concentration inequalities for a permutation-invariant quantum state possibly with an additional symmetry and a sequence of outcomes of independent and identical quantum measurements performed on an adversarial quantum state. For the former case, we used the technique from the representation theory of the permutation group, which is well-known in the community of quantum information theory. With this rather conventional technique, we can upper-bound the permutation-invariant state by i.i.d. quantum states. Compared to the previous de-Finetti-type statement [15, 20, 21], our improvement for this case comes from the fact that one can adaptively choose these i.i.d. states depending on the irreducible subspace of the permutation group.
For the latter case, we found an upper bound on the probability of obtaining each sequence of quantum measurement outcomes with a specific type and then obtained a concentration inequality for the probability of obtaining sequences with the atypical types. For this, we developed, to the best of the author’s knowledge, a completely new technique that uses a geometric structure of the set of probability mass functions. Both bounds we obtained are Sanov-type and are thus expected to be tighter than the conventional bound obtained via Azuma’s inequality. In specific models we discussed in Sec. 3, our bounds even outperform the recently developed refinement of Azuma’s inequality [14].
There remain a few questions on our results. One of the questions is the range of applicability of our results compared to the conventional concentration inequalities such as Azuma’s inequality. As mentioned, our results require either permutation symmetry in quantum system or independent and identical measurements while Azuma’s inequality does not. There may be a few QKD protocols in which neither condition appears to hold such as the differential-phase-shift (DPS) protocol. However, the state-of-the-art finite-size security proof of it decomposes the pulse train into blocks with three pulses to ease the analysis [41], which may then make it possible to apply our results by actively imposing a permutation symmetry to the classical data. Thus, detailed classification of the cases in which one can or cannot apply our results is left to a future work.
Another question is when our concentration inequalities outperform Kato’s inequality [14]. From Fig. 3, it seems that our concentration inequalities can achieve smaller $\epsilon$ more quickly than Azuma’s or Kato’s inequality, and thus the use of our inequalities may be better in the case we need to ensure a small failure probability. However, our concentration inequalities have a dependency either on the dimension of the quantum system (Theorem 1) or the number of the measurement outcomes (Theorem 3) while Azuma’s and Kato’s inequalities do not. Thus, which inequality performs better in a given setup is an open problem.